Solve the system
x+y+z=6
x-y+z=0
x-y-z=-2
find the z value of the solution

Answer by Alan3354(69443)

x+y+z=6
x-y+z=0
x-y-z=-2
----------
Add eqn 1 and 3:
2x = 4
x = 2
------
+y+z = 4 from eqn 1
-y+z = -2 from eqn 2
--------- Add
2z = 2
z = 1
------
y = 3